OpenStudy (anonymous):

Mrs. Noether wants the pool to hold between 1400 ft3 and 3200 ft3 of water and be from 6 to 10 feet deep. Mrs. Noether wants the pool to hold between 1400 ft3 and 3200 ft3 of water and be from 6 to 10 feet deep. The dimensions for the Noether pool that fit the above specifications The amount of pool liner material that will be needed to completely line the inside of the pool (sides and bottom) The amount of water that will be needed to fill the pool if the water needs to be 6 inches below the top of the pool please i really need help!!! :(

OpenStudy (anonymous):

Calculating volume: \[V=l*w*h\] Between 1400 and 3200 is a big difference, so half way between those numbers is 2300ft^3 V = 2300, Between 6 and 10 is 8.5, BUT we can technically use any number between so lets use 8 because its more simple. "The dimensions for the Noether pool that fit the above specifications" \[V=l*w*h\] \[2300=8*w*h\] \[287.5=w*h\] From the instructions you've given me, (i dunno if you left anything out), we can choose any width and length, as long as it equals 287.5 so 143.75 x 143.75 = 287.5 Therefore, her specifications COULD be V=2300, L=8, W=143.75 and W=143.75

OpenStudy (anonymous):

The question is very vague so Im not going to keep answering, I have a feeling you're missing important numbers

OpenStudy (reemii):

does the question mention the length, width and height of the pool?

OpenStudy (anonymous):

omfg yes!!!! thank you guys thats what I needed thank you